OpenStudy (anonymous):

A line has a slope of -2 and contains the points (3, 5) and (4, y). The value of y is

jimthompson5910 (jim_thompson5910):

m = (y2 - y1)/(x2 - x1) -2 = (y2 - y1)/(x2 - x1) -2 = (y - 5)/(4 - 3) now solve for y

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (anonymous):

y=3

jimthompson5910 (jim_thompson5910):

good, I'm getting the same